II. Find the expected value, variance, and standard deviation of the following discrete 

probability distributions. (3 items x 15 points) 

 

1.

T -3 5 7 

P(T) 0.27 0.4 0.33 

 

2.

X 0 1 2 3 4 

P(x=X) 0.21 0.44 0.06 0.11 0.18 

3.

Y 27 30 14 43 

P(Y=y) 0.3 0.4 0.15 0.15


1
Expert's answer
2021-05-07T08:55:59-0400

1.

"\begin{array}{|c|c|c|c|}\hline T& -3& 5& 7\\ \hline P(T=t)& 0.27& 0.4& 0.33\\ \hline\end{array}"


Check the sum of probabilities

"0.27+ 0.4+ 0.33=1"


The expected value is

"\bold{E}[T]= -3\cdot0.27+ 5\cdot0.4+ 7\cdot0.33 =3.5"


The variance is

"\bold{Var}[T]=\bold{E}[T^2]-\bold{E}^2[T]=" "\bigg((-3)^2\cdot0.27+ 5^2\cdot0.4+ 7^2\cdot0.33\bigg)-(3.5)^2 =" "16.35"


The standard deviation is

"\sigma[T]=\sqrt{\bold{Var}[T]}=\sqrt{16.35}\approx4.04"



2.

"\begin{array}{|c|c|c|c|c|c|}\hline X &0& 1& 2& 3& 4 \\ \hline P(X=x) &0.21& 0.44& 0.06& 0.11& 0.18\\ \hline\end{array}"


Check the sum of probabilities

"0.21+ 0.44 +0.06+ 0.11+ 0.18=1"


The expected value is

"\bold{E}[X]=" "0\cdot0.21+1\cdot0.44+2\cdot0.06+3\cdot0.11+4\cdot0.18=" "1.61"


The variance is

"\bold{Var}[X]=" "(0^2\cdot0.21+1^2\cdot0.44+2^2\cdot0.06+3^2\cdot0.11+4^2\cdot0.18)-(1.61)^2\approx1.96"


The standard deviation is

"\sigma[X]=\sqrt{1.96}\approx1.40"



3.

"\begin{array}{|c|c|c|c|c|c|}\hline Y& 27& 30& 14& 43 \\ \hline P(Y=y)& 0.3& 0.4& 0.15& 0.15\\ \hline\end{array}"


Check the sum of probabilities

"0.3 +0.4+ 0.15+ 0.15=1"


The expected value is

"\bold{E}[Y]=27\cdot0.3+ 30\cdot0.4 + 14\cdot0.15 + 43\cdot0.15=" "28.65"


The variance is

"\bold{Var}[Y]=" "(27^2\cdot0.3+ 30^2\cdot0.4 + 14^2\cdot0.15 + 43^2\cdot0.15)-(28.65)^2\approx64.63"


The standard deviation is

"\sigma[Y]=\sqrt{64.63}\approx8.04"